OpenStudy (anonymous):

3. Which of the following equations has an infinite number ofsolutions?

OpenStudy (anonymous):

A: 3x – 3 = –4x B: 2y + 4 – y = 16 C: 7x + 5 = 4x + 5 + 3x D: 6y – 2 = 2(y – 1) plz help??

OpenStudy (gorv):

c

OpenStudy (anonymous):

oh, okay thanks :D

OpenStudy (gorv):

7x+5=4x + 5 + 3x 7x+5=7x+5 it is true for every value of x